OpenStudy (anonymous):

Has anyone proved that y=2y zero on that problem in the first lecture where he used symmetry to get that y=2y zero? Could you help because I keep getting y=2/x zero.

OpenStudy (stacey):

I would need to see the actual problem.

OpenStudy (anonymous):

y-y zero=-1/(x zero)^2 (x-x zero) is the problem. The professor solved for x by plugging in y=0 and got x=2x zero which I understand. Then he said by symmetry you get y=2y zero and we could prove it by taking the same equation and plugging in x=0. Every time I do that, I get y= 2/x zero, not y=2y.

OpenStudy (stacey):

Based on what you are saying, I am assuming that previously in the lecture, it was stated that \[y _{0} = -\frac{ 1 }{ x _{0} }\]If that is the case, then \[x _{0} = -\frac{ 1 }{ y _{0} }\] and \[y=2y _{0}\]

OpenStudy (anonymous):

i must be doing something wrong...thanks anyway

OpenStudy (anonymous):

The area of the triangle as he derives is 2XoYo. As y = f(x) = 1/X, f(Xo) = 1/Xo. Therefore Yo = 1/Xo. Thus, 2XoYo = 2Xo(1/Xo) = 2